a) Calculate velocity of electron in first Bohr orbit of hydrogen atom (Given r = a0). 


 b) Find de-Broglie wavelength of the electron in first Bohr orbit. 


 c) Find the orbital angular momentum of 2p orbital in terms of h/2π units. 


Grade:11

3 Answers

Sunil Kumar FP
9 years ago

a)we have
mvr=nh/2pi
r=a0 =.529A0
v=nh/2pimr
=2.18*10^6 m/s,here for first bohr orbit,n=1
b)debroglie wavelength =h/mv =6.6*10^-34/9.1*10^-31*2.18*10^6
=3.3A0
c)For 2 p value of l=1
Orbital angular momentum=sqrt(l(l+1))h/2pi =sqrt2 h/2pi


Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

a)we have mvr=nh/2pi
r=a0 =.529A0
v=nh/2pimr =2.18*10^6 m/s,
here for first bohr orbit,n=1
b)debroglie wavelength =h/mv
=6.6*10^-34/9.1*10^-31*2.18*10^6
=3.3A0
c)For 2 p value of l=1
Orbital angular momentum = sqrt(l(l+1))h/2pi
= sqrt2 h/2pi

Thanks and Regards
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

a)we have mvr=nh/2pi
r=a0 =.529A0
v=nh/2pimr =2.18*10^6 m/s,
here for first bohr orbit,n=1
b)debroglie wavelength =h/mv
=6.6*10^-34/9.1*10^-31*2.18*10^6
=3.3A0
c)For 2 p value of l=1
Orbital angular momentum = sqrt(l(l+1))h/2pi
= sqrt2 h/2pi

Thanks and Regards